OpenStudy (anonymous):

When it says: Give a constructional definition of the following sets What does this mean?

OpenStudy (fibonaccichick666):

I need context, (what sets)

OpenStudy (anonymous):

sorry was afk @FibonacciChick666

OpenStudy (fibonaccichick666):

I'm about to be afk myself I have class

OpenStudy (anonymous):

oh sorry im from the UK lmao

OpenStudy (anonymous):

I'm going to bed but ill be on when i wake so if you could answer it i would appreciate it @fibonaccichick666

OpenStudy (fibonaccichick666):

So the issue I am having is that what you have linked to, is what I know as a constructional definition. @ganeshie8 , any ideas on another interpretation?

ganeshie8 (ganeshie8):

I think it gives all powers of 2

OpenStudy (fibonaccichick666):

no, it would include 9. Right?

ganeshie8 (ganeshie8):

\(3 | 9\) so 9 is out

OpenStudy (fibonaccichick666):

ah division, I took it as n cannot be a prime

OpenStudy (fibonaccichick666):

hmm, well, Does that mean the prime does not divide itself as well? If not, then I agree

ganeshie8 (ganeshie8):

does this look okay \[\mathbb{ \{ n \in Z : n=2^k \text{for some } k \ge 0\} } \]

OpenStudy (fibonaccichick666):

Won't that end in a zero at some point?

ganeshie8 (ganeshie8):

2 4 8 6 2 ...

OpenStudy (fibonaccichick666):

ok, how about the ending in sixes, could they be divided by 3?

ganeshie8 (ganeshie8):

2^k is never divisible by 3 cuz there is no 3 in its prime factorization ;p

OpenStudy (fibonaccichick666):

I guess, can that be guaranteed for every k?

ganeshie8 (ganeshie8):

i think you're asking about proving fundametal theorem of arithmetic

OpenStudy (fibonaccichick666):

oh gosh, no you're right

OpenStudy (fibonaccichick666):

I was just thinking in terms of the last digit

ganeshie8 (ganeshie8):

that theorem guaranteees that the prime factorization is unique upto permutations.. so the prime power like : 2^k wont split into 3^m*5^n or some other thing ever...

OpenStudy (fibonaccichick666):

but, for this we can say that by the (theorem that states all numbers have a prime factorization) we can express every number as \[n=p_1p_2p_3.....p_k\] by setting the limitation of n not being allowed to be divisible by any prime greater than or equal to 3, we are limited in our prime choices leading us to any n must be able to be written as \[2^k\]

ganeshie8 (ganeshie8):

thats a nice argument

OpenStudy (fibonaccichick666):

It was your catch that made it possible :) I have an issue with reading division backwards when expressed with |, so I was like, well that's the definition of a prime set that includes 1 and 2,